### v1

@@ 5 CONCLUSIONS
We proposed the first mass conservation study for 3D printing simulations based on the conservative level set method. There is no clear guide on how to tune the level-set parameters, and their effects on the cross-sectional area of a printed strand have not been quantified before. The numerical results indicate that by carefully reducing the reinitialization parameter and the interface thickness, the errors on the cross-sectional area are reduced. This indicates a better enforcement of mass conservation in the simulation. Although decreasing the level-set parameters seems logical, we found that the reductions in error tend to decrease at some point, and that excessive reduction can lead to higher computational cost. Furthermore, we also found that the requirements on mesh refinements that are common for the level-set method can be attenuated by selecting an adequate interface thickness. Normally, the recommendation in commercial software is to tune the reinitialization parameter, but the results show that careful refinement of the interface thickness can have a more significant impact on the mass conservation without compromising the computational cost. Finally, the best level set parameters found were used to simulate 3D printing scenarios with different settings. The quantitative results had small errors, and the cross-sectional shapes agree well with the experimentally validated results in the literature.
@@ 6 ACKNOWLEDGMENTS
The authors would like to thank Tusher Mollah and Monica Zakhari for the valuable discussions about the simulation settings and initial results. Furthermore, the authors would like to acknowledge the financial support from NWO (grant number: 18763), Wageningen Food & Biobased Research, TNO, Ruitenberg Ingredients B.V., LambWeston Holdings, Inc., and Gastronology 3D Food Works.

### v2

@@ 5 CONCLUSIONS
This work presented a systematic mass conservation analysis of MEX-AM simulations using the conservative level-set method. While the CLS formulation is attractive due to its smooth interface representation and numerical robustness, clear guidelines for selecting its key parameters have been lacking. By directly relating mass conservation errors to deviations in the steady-state cross-sectional area of printed strands, we provided a quantitative framework to assess the impact of the reinitialization parameter and interface thickness.
For single-layer deposition, the results show that mass conservation errors strongly depend on both the reinitialization parameter ($\gamma$) and the interface thickness ($\epsilon$). The results demonstrate that proper tuning of the parameters decreased mass conservation errors, leading to more accurate predictions of strand geometry. However, excessive reduction of the level-set parameters yielded diminishing improvements while increasing computational cost. Among the two parameters, the interface thickness was found to have a stronger influence on mass conservation accuracy than the reinitialization parameter, reducing the error below $1\%$. Proper selection of the interface thickness not only reduces steady-state errors but also enables the use of coarser meshes without significant compromises in accuracy, thereby alleviating the stringent mesh requirements and computational cost typically associated with level-set-based simulations. This finding highlights that interface thickness selection can be more effective than aggressive mesh refinement or reinitialization tuning in improving CLS-based simulation accuracy.
The effectiveness of the proposed parameter selection strategy was further validated for different printing speeds and nozzle-to-bed distances. For all tested cases of single-layer printing, mass conservation errors remained below $2\%$ Furthermore, the methodology was also successfully extended to multilayer simulations involving Bingham-type fluids and different nozzle geometries. In two-layer concrete and metakaolin printing simulations, the total cross-sectional area errors were below $0.5\%$ for concrete and below $3\%$ for metakaolin, while the predicted strand shapes showed good qualitative and quantitative agreement with experimentally validated results available in the literature.
Despite these encouraging results, the inherent limitations of the CLS formulation must be acknowledged. Due to the use of a diffuse interface with finite thickness and smooth Heaviside regularization of material properties, deposited layers are not preserved as sharp, distinct interfaces. In multilayer simulations, previously formed interfaces are smoothed, and there is partial blending between layers. While using a diffuse interface improves numerical stability and avoids spurious oscillations, it limits the ability of the CLS method to accurately represent sharp interlayer boundaries or material discontinuities that may be relevant for mechanical performance and interlayer bonding analysis.
Future work will focus on incorporating adaptive mesh refinement focused on the evolving interface, extending the analysis to viscoelastic and thixotropic material models, as well as integrating curing or solidification mechanisms, to further improve the predictive capability of the simulations. Finally, the quantitative mass conservation metrics developed in this work provide practical and transferable guidelines for tuning conservative level-set parameters in MEX-AM simulations and could be leveraged for automated parameter tuning.
@@ 6 ACKNOWLEDGMENTS
The authors would like to thank Monica Zakhari for the valuable discussion about the simulation settings and initial results. Furthermore, the authors would like to acknowledge the financial support from NWO (grant number: 18763), Wageningen Food & Biobased Research, TNO, Ruitenberg Ingredients B.V., LambWeston Holdings, Inc., and Gastronology 3D Food Works.